OpenStudy (thomas5267):

I have taken a screening test today on a summer course for Linear Algebra. I would like to have the solution for question 4 and some hints on question 10. Furthermore, I solved question 8 with angle sum identity but would like to know a way that wouldn't require it.

OpenStudy (thomas5267):

Here is the question booklet.

OpenStudy (anonymous):

alpha and beta are the roots ,right?

OpenStudy (thomas5267):

Yep. I am thinking something like \[(-x+\alpha)(x-\beta)\] but couldn't finish it.

OpenStudy (anonymous):

i dont think it is solved that way now my memory is bit rusty,i solved these like 8 years ago? i will google and fill in my blank spaces

OpenStudy (anonymous):

the y intercept is 0,4 then one of the x's value is 4

OpenStudy (anonymous):

when f(x) = 0

OpenStudy (thomas5267):

c is 4 but I couldn't figure the rest out.

OpenStudy (anonymous):

how is c 4 O.o

OpenStudy (thomas5267):

\[ \begin{align*} F(x)&=-x^2+bx+c\\ 4&=-0^2+0b+c\\ c&=4 \end{align*} \]

OpenStudy (anonymous):

hey,no... one of x's values is 4,not whole function's value

OpenStudy (thomas5267):

Isn't the y-intercept of F(X) (0,4)?

OpenStudy (anonymous):

yes

OpenStudy (thomas5267):

Doesn't F(0)=4 implies c=4? (or am I mad today?)

OpenStudy (anonymous):

*shrugs* okay am confused with these basics only wait @ganeshie8 how come he took f(0) = 4?

OpenStudy (thomas5267):

Isn't (0,4) a point on F(x)?

ganeshie8 (ganeshie8):

yes so c = 4

ganeshie8 (ganeshie8):

\( F(x) =-x^2+bx+4 = -(x-\alpha )(x-\beta ) \)

ganeshie8 (ganeshie8):

comparing x coefficients and constants terms both sides you get : \(\alpha + \beta = b\) \(\alpha \beta = -4\)

ganeshie8 (ganeshie8):

and you're given another equation : \( \beta-\alpha = 4\)

ganeshie8 (ganeshie8):

3 equations, 3 unknowns - you can solve

OpenStudy (anonymous):

thanks @ganeshie8 gladly i didnt screw him with my half knowledge abt this

OpenStudy (thomas5267):

Practising LaTeX \[ \begin{align*} \alpha\beta&=-4\\ \beta&=-\frac{4}{\alpha} \end{align*} \\ \] \[ \begin{align*} \beta-\alpha&=4\\ -\frac{4}{\alpha}-\alpha&=4\\ -4-\alpha^2&=4\alpha\\ \alpha^2+4\alpha+4&=0\\ (\alpha+2)^2&=0\\ \alpha&=-2\\ \beta&=2\\ b&=0 \end{align*} \\ \]

ganeshie8 (ganeshie8):

haha looks great !! ive never learned these alignment stuff before lol xD

OpenStudy (thomas5267):

Is it possible to deduce that (0,4) is the vertex given that \(F(x)=-x^2+bx+4\)

ganeshie8 (ganeshie8):

definitely, now that u have b value, plug it in and find the vertex

OpenStudy (thomas5267):

I mean without solving for b. If we can deduce that (0,4) is the vertex, then the roots would clearly be (-2,0) and (2,0) as the x-coordinate of the vertex must lie between the midpoint of the x-coordinates of the roots (by the symmetry of parabola and \(\beta-\alpha=4\)).

OpenStudy (thomas5267):

Just trying to find shortcuts lol.

OpenStudy (thomas5267):

I will close this question and repost this so that people can earn more medals. After all, 3 questions and only 1 best medal seems unfair.

ganeshie8 (ganeshie8):

il try this later again :)